1.) If the marginal propensity to consume is 0.75 and investment spending increases by $200 billion,...

1.)

If the marginal propensity to consume is 0.75 and investment spending increases by $200 billion, equilibrium GDP will increase by____.

$350 billion

$150 billion

$200 billion

$266.7 billion

$800 billion

2.)

AE = 3000 + 0.75*RGDP. Given this equation for AE, find equilibrium GDP

$1,000

$750

$12,000

$2,250

3.)

The four components of aggregate planned expenditure are

the real interest rate, disposable income, wealth, and expected future income

the real interest rate, consumption expenditure, investment, and government expenditures

consumption expenditure, investment, government expenditures, and wealth

consumption expenditure, investment, government expenditures, and net exports

4.)

When disposable income decreases

the consumption function shifts up

there is a rightward movement along the consumption function

the consumption function shifts down

there is a leftward movement along the consumption function
Answer #1

1) Solution: $800 billion

Explanation: Multiplier = 1 / (1-MPC) = 1 / (1-0.75) = 4

Increase in equilibrium GDP = 4 * $200 billion = $800 billion

 

2) Solution: $12,000

Explanation: AE = 3000 + 0.75*RGDP. Given this equation for AE, find equilibrium GDP

Here Multiplier = 1 / (1-MPC) = 1 / (1-0.75) = 4

Thus equilibrium GDP = 3000 * 4 = 12,000

 

3) Solution: consumption expenditure, investment, government expenditures, and net exports

Explanation: The four components of planned aggregate expenditures includes investment, consumption, government expenditures, and net exports

 

4) Solution: there will be a leftward movement along the consumption function

Explanation: The changes in disposable income may cause movements along the consumption function